At a basketball game, a vender sold a combined total of 175 sodas and hot dogs. The number of sodas sold was 55 more than the number of hot dogs sold. Find the number of sodas sold and the number of hot dogs sold.
Answer by ikleyn(52779)

.

x = # of hot dogs

(x+55) = # of sodas.


Combines

    x + (x+55) = 175.


Simplify and find x

    2x = 175 - 55

    2x =    120

     x =    120/2 = 60.


ANSWER.  60 hot dogs and 60+55 = 115 sodas.


CHECK.   60 + 115 = 175 sold combined.    ! Correct !

Solved.